OpenStudy (anonymous):

What is the formula for the square of a sum? a.)2a2 + ab - b2 b.)a2 + ab + 2b2 c.)a2 + 2ab + b2 d.)a2 - 2ab + b2

OpenStudy (anonymous):

try it for yourself (a+b) * (a+b) this is where you have to use that FOIL thing

OpenStudy (anonymous):

you end up with a*a + a*b + b*a + b*b the a*b and b*a are the same thing, so when you add them they become 2ab a*a = a^2 b*b = b^2 so you get a^2 + 2ab + b^2